OpenStudy (anonymous):

I have to find the volume of a cone. the radius is 4 and the height is 3. i complete half the problem but i am unsure what to do with that 1/3. so i have v=1\3π14=48. need help

OpenStudy (anonymous):

\(v=\frac{1}{3}\pi r^2h\) \(v=\frac{1}{3}\pi(4^2)(3)\) \(v=\frac{1}{3}\pi48\)

OpenStudy (anonymous):

From there, you multiply the 1/3 by 48 or divide 48 by 3.

OpenStudy (anonymous):

That will give you the answer in terms of \(\pi\).